# Few-Shot-Based Modular Image-to-Video Adapter for Diffusion Models

### v1

@@ 2.1 Image animation (I2V) DMs
Image animation is a video generation task aiming to produce realistic motion from a static image input, often one of the final video frames. An intuitive idea is to impose an image-conditioning mechanism on a pre-trained T2V model, but the design of such mechanisms is nontrivial. Options vary from 1) integrating the image into the noise signal via forward diffusion process [cil], concatenation channel-wise [pia] or frame-wise [lamp, $a$, consisti2v, trip, cinemo], 2) imposing classifier-free guidance (CFG)[cfg] with a feature encoding the image [svd, ip-adapter, videocrafter1] (e.g. by a CLIP image encoder [clip]), and 3) doing both [videocomposer, dynamicrafter, i2v-adapter, i2vgen]. It is disputable whether image-embedding-based CFG can benefit I2V generation performance. [animatelcm, cil, cinemo, consisti2v] criticize that CLIP image encoder focuses more on the high-level semantic information rather than the visual details, which are more important in the I2V problem. [dynamicrafter, ip-adapter] propose to learn a projection module upon the CLIP image embedding to introduce inductive bias towards visual details, but this requires training on large-scale datasets. [cinemo, trip] guide the DM to learn the difference between each frame and the reference image to enhance temporal coherency, and some methods take additional control signals from users, such as the mask of the subject(s) to animate [$a$] and motion trajectories [mofa, motioni2v] , which are data-demanding as well. We note some methods specializing in specific motion types, such as the oscillating motion of foreground objects [gid], motion of hair [hair], portrait [liveportrait, xportrait] and body [animateanyone, dreampose] etc., while this paper is concerned with open-domain motion patterns. We also highlight a related task known as motion transfer [flexiact], which involves generating a video of a specific subject mimicking the motion in a reference video. Unlike the general-purpose I2V problem studied in this paper, which is designed to generalize across different subjects, motion transfer typically necessitates subject-specific fine-tuning, limiting its scalability and adaptability.
@@ 3.2 Problem Setting and Motivation
We adopt the problem setting introduced by LAMP [lamp], which trains a DM to animate an input image II with a target motion pattern $\mathcal{A}$, using a small-scale dataset (size at $10^{1}$ scale) specific to the motion pattern. The output is a video $V=\{I^{i}|i=1,\dots,F\}$ with $I^{1}\equiv I$ serving as the initial frame. While we build upon this setup, our work introduces two key modifications aimed at improving training stability and controllability, as detailed below.
@@ 3.3.1 Cross-frame Attention (CFA) Layers
Previous studies treat $\{\lambda_{1,2,3}\}$ as fixed hyperparameters. Options include all 11s in Tune-A-Video, $(0.7,0.15,0.15)$ in PoseAnimate [poseanimate] and $(1,1,0)$ in I2V-Adapter [i2v-adapter]. However, we find the weighting of CFA layers non-trivial due the diversity of motion patterns. In general, faster or erratic motion (e.g., raindrops) requires higher $\lambda_{2}$ to enforce consistency with $I^{1}$, which anchors global appearance. Meanwhile, slower or subtler motion (e.g., humans, animals) favors a higher $\lambda_{3}$ to maintain continuity with the previous frame. The $\lambda$-assignment problem is even more intricate when considering the diffusion time step: given the insight that different visual contents are formed at different time steps [directed], the optimal $\lambda$ configuration can shift over time.
To alleviate the burden of manual $\lambda$-assignment, we propose an adaptive weighting module $\phi$, attached to each SA layer. This module is defined as $\phi(\mathbf{c}_{t};\mathbf{W}_{\phi})=\sigma(\mathrm{SiLU}(\mathbf{c}_{t})\mathbf{W}_{\phi})$ where $\mathbf{c}_{t}\in\real^{d_{t}}$ is the time step embedding and $\mathbf{W}_{\phi}\in\real^{d_{t}\times 2}$ is a learnable projection. This module dynamically outputs a 2-dimensional weighting vector, which is assigned to $(\lambda_{2},\lambda_{3})$, while $\lambda_{1}$ is fixed to 11. We demonstrate the impact of $\phi$ in the ablation study section.
@@ 3.3.2 CA Layers with Implicit Prompt
Although our modular I2V framework does not rely on text prompts from users, we can still leverage the priors within the CA layers of the base model. Since each MIVA specializes in a single motion type, we associate it with a fixed prompt that semantically represents its target motion pattern. We also assign this prompt uniformly to all training videos exhibiting the target motion. Appearance information is omitted from the prompt, as it is already provided by II. Rather than using natural language, we encode this motion-specific prompt as a learnable embedding $\mathbf{c}$, which we refer to as an implicit prompt, eliminating the need for manual prompt engineering.
@@ 3.4 Masked MIVA (M-MIVA)
The lightweight design of MIVA enables training in few-shot scenarios without requiring manual annotations. However, in the absence of constraints, MIVA may indiscriminately learn from all moving pixels in the training samples. This makes it susceptible to data imperfections, such as incidental motion from unrelated subjects, and biases inherent in limited samples, often leading to visible artifacts (see Fig. 8). To improve robustness against data imperfections, we incorporate additional cues that help the DM focus on the subject of interest. Inspired by Through-The-Mask [through-the-mask], we utilize subject-specific mask sequences to distinguish foreground from background within each frame, effectively guiding the model’s attention. We employ an off-the-shelf detect-and-segment pipeline [groundingdino, sam2], which requires no additional user input, to extract the subject masks from sample videos during training and from II during inference.
@@
As illustrated in Fig. 3, M-MIVA operates on a video tensor $\mathbf{x}^{1:F};\mathbf{s}^{1:F}]$, formed by concatenating the generated frame sequence $\{\mathbf{x}^{1:F}\}$ and its corresponding mask sequence $\{\mathbf{s}^{1:F}\}$. The initial frame and mask, $\mathbf{x}^{1}$ and $\mathbf{s}^{1}$, are set to II and its subject mask in the latent domain. M-MIVA can be trained using the same denoising objective as the base DM, with key modifications to the attention layers that enable interaction between video and mask modalities. Specifically, we introduce a parallel mask-modality attention stream to each SA layer, composed of independently parameterized CFA layers. Each frame-mask CFA pair is identically initialized to preserve initial symmetry. In addition, the mask sequence is transformed into attention masks that guide the attention computation in the video-modality stream, as described in the following subsection.
@@ 4.3.1 CFA-Weighting Modules $\phi$
We observe substantial variation in the CFA weights $\{\lambda_{2,3}\}$ across layers, time steps, and motion types. We select the bottleneck block of the U-Net, of lowest resolution and thus conveying most semantic information, to shed a light on the association between $\lambda_{2}$ and the motion pattern. A larger $\lambda_{2}$ indicates stronger dependency on the input image II and, on the other hand, less dependency on the previous frame. As shown in Fig. 6, the rain and waterfall cases exhibit the highest $\lambda_{2}$, indicating minimal reliance on the previous frame due to the complex, fast-evolving textures of fluid dynamics. Conversely, $\lambda_{2}$ are lowest for horse and human motion, consistent with the expectation that body and facial animation require greater temporal coherency, wherein the previous frame plays a more important role.
As shown in Fig. 7, we conduct a comparison study to assess the impact of $\phi$, where MIVAs with $\lambda_{2}\equiv\lambda_{3}\equiv 0.5$ serve as the baseline. In case $\mathcal{R}$, the adaptive mechanism assigns larger $\lambda_{2}$, emphasizing the input frame II and promoting better preservation of visual details, while the baseline disproportionately attends to raindrops, resulting in background blur. Meanwhile, case $\mathcal{H}$ poses a greater challenge due to subtle limb movements; existing approaches and the baseline often produce anatomically inconsistent outputs with unrealistic leg counts or body structure. By biasing the CFA weights toward the previous frame, the adaptive mechanism enhances motion continuity and significantly improves anatomical fidelity in maintaining the realism of body appearance, especially that of the legs. Quantitatively, the introduction of $\phi$ yields consistent gains across most objective evaluation metrics, as reported in Tab. 1.
@@ B.2.1 Parallelizing M-MIVAs
where $\epsilon$ is a small positive constant added for numerical stability, and $\mathbf{1}(x)$ denotes an indicator function that returns 11 if $x$ is true and 0 otherwise. $M^{*}$ is then applied to modulate the SA layers.
@@ B.3.1 Pre-processing Techniques
Pre-processing techniques manipulate the initial noise at inference time $\mathbf{x}_{T}$ to replace the original white Gaussian noise. The underlying intuition is that $\mathbf{x}_{T}$ should preserve a small amount of information from II to align with the forward process, rather than being left completely random [cil]. We discover that integrating the shared-noise mechanism by LAMP [lamp] and the DCTInit algorithm by Cinemo [cinemo] results in more desirable frame quality and fidelity to II compared to applying either method alone. We summarize the pre-processing pipeline in Alg. 1.
@@ B.3.2 Post-processing Techniques
Post-processing techniques manipulate the output of DM, either during the sampling process (e.g., Adaptive Instance Normalization or AdaIN [adain]) or after sampling (e.g., histogram matching with II). These techniques are also tuning-free. In contrast to LAMP which applies AdaIN after each sampling iteration, we observe that applying AdaIN only at the end of the last iteration, right before the VAE decoding step, yields consistently higher video frame quality. We discard the histogram matching step in LAMP, arguing that an equivalent histogram with II is a poor proxy for temporal coherency, effective only in cases of little dynamics.
@@ B.4.2 MIVA based on AnimateDiff (U-Net)
Training follows the standard DM training objective, minimizing the denoising loss. We use the Adam optimizer [adam] with a fixed learning rate of $10^{-5}$, excluding the first frame from the loss computation. Training is conducted over 15K iterations on a single GPU, with a minibatch size of 1. The VRAM usage is 9.8 GB for MIVA and 11.0 GB for M-MIVA. At inference time, we employ DDIM sampling [ddim] with 50 time steps. The total number of learnable parameters in MIVA is 35.4M (24.8M in CFA layers, 1.6M in CA layers, and 9.0M in t-SA layers). This corresponds to approximately 3% of the base model’s 1.29B parameters. M-MIVA introduces an additional 20.4M learnable parameters for the mask generation stream, increasing the total to roughly 5%. We set the classifier-free guidance (CFG [cfg]) scale to 11 to avoid training and inference of unconditional generation.
@@ B.5.1 IP-Adapter
We follow IP-Adapter by injecting an additional attention operation term for each CA layer, with $K$ and $V$ based on the embedding of II, given by the CLIP image encoder [clip] further processed by a pre-trained IP-Adapter. Like I2V-Adapter, we train the learnable adapter weights while freezing the IP-Adapter weights. However, we observe this harms the resulting quality drastically: as shown in Fig. S11, the generated videos exhibit severe blurring and desaturation artifacts. A possible reason is that the new image-conditioned term may distort the distribution of the CA layer output, requiring more data for the learnable weights to adapt.
@@ C.2.5 Aesthetic Quality Loss
concerns the aesthetic appeal of each frame measured by the LAION aesthetic predictor [laion]. Each frame receives a rating between 0 and $10$, which is later linearly normalized to $[0,100]$. Considering that input images vary in perceptual quality, we report the loss of quality compared to II rather than the original scores. The final score for a video is obtained by averaging the quality loss of all frames.
@@ D.1 Adjusting MIVA Weights in Multi-motion-pattern Animation
Section 2-2 in the Project Page showcases how adjusting the MIVA weights enables fine-grained control over the intensity of each constituent motion pattern. Empirically, a total weight of 11 yields the most stable animation results.

### v2

@@ 2.1 Image animation (I2V) DMs
Image animation is a video generation task aiming to produce realistic motion from a static image input, often one of the final video frames. An intuitive idea is to impose an image-conditioning mechanism on a pre-trained T2V model, but the design of such mechanisms is nontrivial. Options vary from 1) integrating the image into the noise signal via forward diffusion process [cil], concatenation channel-wise [pia] or frame-wise [lamp, aa, consisti2v, trip, cinemo], 2) imposing classifier-free guidance (CFG)[cfg] with a feature encoding the image [svd, ip-adapter, videocrafter1] (e.g. by a CLIP image encoder [clip]), and 3) doing both [videocomposer, dynamicrafter, i2v-adapter, i2vgen]. It is disputable whether image-embedding-based CFG can benefit I2V generation performance. [animatelcm, cil, cinemo, consisti2v] criticize that CLIP image encoder focuses more on the high-level semantic information rather than the visual details, which are more important in the I2V problem. [dynamicrafter, ip-adapter] propose to learn a projection module upon the CLIP image embedding to introduce inductive bias towards visual details, but this requires training on large-scale datasets. [cinemo, trip] guide the DM to learn the difference between each frame and the reference image to enhance temporal coherency, and some methods take additional control signals from users, such as the mask of the subject(s) to animate [aa] and motion trajectories [mofa, motioni2v] , which are data-demanding as well. We note some methods specializing in specific motion types, such as the oscillating motion of foreground objects [gid], motion of hair [hair], portrait [liveportrait, xportrait] and body [animateanyone, dreampose] etc., while this paper is concerned with open-domain motion patterns. We also highlight a related task known as motion transfer [flexiact], which involves generating a video of a specific subject mimicking the motion in a reference video. Unlike the general-purpose I2V problem studied in this paper, which is designed to generalize across different subjects, motion transfer typically necessitates subject-specific fine-tuning, limiting its scalability and adaptability.
@@ 3.2 Problem Setting and Motivation
We adopt the problem setting introduced by LAMP [lamp], which trains a DM to animate an input image $I$ with a target motion pattern $\mathcal{A}$, using a small-scale dataset (size at $10^{1}$ scale) specific to the motion pattern. The output is a video $V=\{I^{i}|i=1,\dots,F\}$ with $I^{1}\equiv I$ serving as the initial frame. While we build upon this setup, our work introduces two key modifications aimed at improving training stability and controllability, as detailed below.
@@ 3.3.1 Cross-frame Attention (CFA) Layers
Previous studies treat $\{\lambda_{1,2,3}\}$ as fixed hyperparameters. Options include all $1$s in Tune-A-Video, $(0.7,0.15,0.15)$ in PoseAnimate [poseanimate] and $(1,1,0)$ in I2V-Adapter [i2v-adapter]. However, we find the weighting of CFA layers non-trivial due the diversity of motion patterns. In general, faster or erratic motion (e.g., raindrops) requires higher $\lambda_{2}$ to enforce consistency with $I^{1}$, which anchors global appearance. Meanwhile, slower or subtler motion (e.g., humans, animals) favors a higher $\lambda_{3}$ to maintain continuity with the previous frame. The $\lambda$-assignment problem is even more intricate when considering the diffusion time step: given the insight that different visual contents are formed at different time steps [directed], the optimal $\lambda$ configuration can shift over time.
To alleviate the burden of manual $\lambda$-assignment, we propose an adaptive weighting module $\phi$, attached to each SA layer. This module is defined as $\phi(\mathbf{c}_{t};\mathbf{W}_{\phi})=\sigma(\mathrm{SiLU}(\mathbf{c}_{t})\mathbf{W}_{\phi})$ where $\mathbf{c}_{t}\in\real^{d_{t}}$ is the time step embedding and $\mathbf{W}_{\phi}\in\real^{d_{t}\times 2}$ is a learnable projection. This module dynamically outputs a 2-dimensional weighting vector, which is assigned to $(\lambda_{2},\lambda_{3})$, while $\lambda_{1}$ is fixed to $1$. We demonstrate the impact of $\phi$ in the ablation study section.
@@ 3.3.2 CA Layers with Implicit Prompt
Although our modular I2V framework does not rely on text prompts from users, we can still leverage the priors within the CA layers of the base model. Since each MIVA specializes in a single motion type, we associate it with a fixed prompt that semantically represents its target motion pattern. We also assign this prompt uniformly to all training videos exhibiting the target motion. Appearance information is omitted from the prompt, as it is already provided by $I$. Rather than using natural language, we encode this motion-specific prompt as a learnable embedding $\mathbf{c}$, which we refer to as an implicit prompt, eliminating the need for manual prompt engineering.
@@ 3.4 Masked MIVA (M-MIVA)
The lightweight design of MIVA enables training in few-shot scenarios without requiring manual annotations. However, in the absence of constraints, MIVA may indiscriminately learn from all moving pixels in the training samples. This makes it susceptible to data imperfections, such as incidental motion from unrelated subjects, and biases inherent in limited samples, often leading to visible artifacts (see Fig. 8). To improve robustness against data imperfections, we incorporate additional cues that help the DM focus on the subject of interest. Inspired by Through-The-Mask [through-the-mask], we utilize subject-specific mask sequences to distinguish foreground from background within each frame, effectively guiding the model’s attention. We employ an off-the-shelf detect-and-segment pipeline [groundingdino, sam2], which requires no additional user input, to extract the subject masks from sample videos during training and from $I$ during inference.
@@
As illustrated in Fig. 3, M-MIVA operates on a video tensor $\mathbf{x}^{1:F};\mathbf{s}^{1:F}]$, formed by concatenating the generated frame sequence $\{\mathbf{x}^{1:F}\}$ and its corresponding mask sequence $\{\mathbf{s}^{1:F}\}$. The initial frame and mask, $\mathbf{x}^{1}$ and $\mathbf{s}^{1}$, are set to $I$ and its subject mask in the latent domain. M-MIVA can be trained using the same denoising objective as the base DM, with key modifications to the attention layers that enable interaction between video and mask modalities. Specifically, we introduce a parallel mask-modality attention stream to each SA layer, composed of independently parameterized CFA layers. Each frame-mask CFA pair is identically initialized to preserve initial symmetry. In addition, the mask sequence is transformed into attention masks that guide the attention computation in the video-modality stream, as described in the following subsection.
@@ 4.3.1 CFA-Weighting Modules $\phi$
We observe substantial variation in the CFA weights $\{\lambda_{2,3}\}$ across layers, time steps, and motion types. We select the bottleneck block of the U-Net, of lowest resolution and thus conveying most semantic information, to shed a light on the association between $\lambda_{2}$ and the motion pattern. A larger $\lambda_{2}$ indicates stronger dependency on the input image $I$ and, on the other hand, less dependency on the previous frame. As shown in Fig. 6, the rain and waterfall cases exhibit the highest $\lambda_{2}$, indicating minimal reliance on the previous frame due to the complex, fast-evolving textures of fluid dynamics. Conversely, $\lambda_{2}$ are lowest for horse and human motion, consistent with the expectation that body and facial animation require greater temporal coherency, wherein the previous frame plays a more important role.
As shown in Fig. 7, we conduct a comparison study to assess the impact of $\phi$, where MIVAs with $\lambda_{2}\equiv\lambda_{3}\equiv 0.5$ serve as the baseline. In case $\mathcal{R}$, the adaptive mechanism assigns larger $\lambda_{2}$, emphasizing the input frame $I$ and promoting better preservation of visual details, while the baseline disproportionately attends to raindrops, resulting in background blur. Meanwhile, case $\mathcal{H}$ poses a greater challenge due to subtle limb movements; existing approaches and the baseline often produce anatomically inconsistent outputs with unrealistic leg counts or body structure. By biasing the CFA weights toward the previous frame, the adaptive mechanism enhances motion continuity and significantly improves anatomical fidelity in maintaining the realism of body appearance, especially that of the legs. Quantitatively, the introduction of $\phi$ yields consistent gains across most objective evaluation metrics, as reported in Tab. 1.
@@ B.2.1 Parallelizing M-MIVAs
where $\epsilon$ is a small positive constant added for numerical stability, and $\mathbf{1}(x)$ denotes an indicator function that returns $1$ if $x$ is true and 0 otherwise. $M^{*}$ is then applied to modulate the SA layers.
@@ B.3.1 Pre-processing Techniques
Pre-processing techniques manipulate the initial noise at inference time $\mathbf{x}_{T}$ to replace the original white Gaussian noise. The underlying intuition is that $\mathbf{x}_{T}$ should preserve a small amount of information from $I$ to align with the forward process, rather than being left completely random [cil]. We discover that integrating the shared-noise mechanism by LAMP [lamp] and the DCTInit algorithm by Cinemo [cinemo] results in more desirable frame quality and fidelity to $I$ compared to applying either method alone. We summarize the pre-processing pipeline in Alg. 1.
@@ B.3.2 Post-processing Techniques
Post-processing techniques manipulate the output of DM, either during the sampling process (e.g., Adaptive Instance Normalization or AdaIN [adain]) or after sampling (e.g., histogram matching with $I$). These techniques are also tuning-free. In contrast to LAMP which applies AdaIN after each sampling iteration, we observe that applying AdaIN only at the end of the last iteration, right before the VAE decoding step, yields consistently higher video frame quality. We discard the histogram matching step in LAMP, arguing that an equivalent histogram with $I$ is a poor proxy for temporal coherency, effective only in cases of little dynamics.
@@ B.4.2 MIVA based on AnimateDiff (U-Net)
Training follows the standard DM training objective, minimizing the denoising loss. We use the Adam optimizer [adam] with a fixed learning rate of $10^{-5}$, excluding the first frame from the loss computation. Training is conducted over 15K iterations on a single GPU, with a minibatch size of 1. The VRAM usage is 9.8 GB for MIVA and 11.0 GB for M-MIVA. At inference time, we employ DDIM sampling [ddim] with 50 time steps. The total number of learnable parameters in MIVA is 35.4M (24.8M in CFA layers, 1.6M in CA layers, and 9.0M in t-SA layers). This corresponds to approximately 3% of the base model’s 1.29B parameters. M-MIVA introduces an additional 20.4M learnable parameters for the mask generation stream, increasing the total to roughly 5%. We set the classifier-free guidance (CFG [cfg]) scale to $1$ to avoid training and inference of unconditional generation.
@@ B.5.1 IP-Adapter
We follow IP-Adapter by injecting an additional attention operation term for each CA layer, with $K$ and $V$ based on the embedding of $I$, given by the CLIP image encoder [clip] further processed by a pre-trained IP-Adapter. Like I2V-Adapter, we train the learnable adapter weights while freezing the IP-Adapter weights. However, we observe this harms the resulting quality drastically: as shown in Fig. S11, the generated videos exhibit severe blurring and desaturation artifacts. A possible reason is that the new image-conditioned term may distort the distribution of the CA layer output, requiring more data for the learnable weights to adapt.
@@ C.2.5 Aesthetic Quality Loss
concerns the aesthetic appeal of each frame measured by the LAION aesthetic predictor [laion]. Each frame receives a rating between 0 and $10$, which is later linearly normalized to $[0,100]$. Considering that input images vary in perceptual quality, we report the loss of quality compared to $I$ rather than the original scores. The final score for a video is obtained by averaging the quality loss of all frames.
@@ D.1 Adjusting MIVA Weights in Multi-motion-pattern Animation
Section 2-2 in the Project Page showcases how adjusting the MIVA weights enables fine-grained control over the intensity of each constituent motion pattern. Empirically, a total weight of $1$ yields the most stable animation results.